# Bidding in Ancillary Service Markets: An Analytical Approach Using Extreme Value Theory

### v2

@@ Abstract
To encourage the participation of stochastic distributed energy resources in Nordic ancillary service markets, the Danish transmission system operator, Energinet, has introduced grid codes requiring a minimum 90% reliability for the full availability of reserve capacity bids. This paper addresses the bidding strategy of flexibility aggregators under Energinet’s reliability requirement by proposing a chance-constrained optimization model. An analytical solution is developed using ideas from extreme value theory, focusing on the “tail” of the empirical data used for flexibility estimation, capturing extreme events where failures are more likely to occur. The proposed model is applied to an electric vehicle aggregator participating in the Nordic market for frequency containment reserve for disturbances (FCR-D). Our results from a realistic case study show that the proposed analytical solution outperforms a commonly used sample-based approach in terms of out-of-sample constraint violation rate.
Index Terms:
Stochastic flexibility, ancillary services, bidding strategy, chance-constrained optimization, extreme value theory
I Introduction
I-A Background and motivation
Flexibility aggregators can pool and manage smaller flexible units located on the consumer side of power systems, known as distributed energy resources. These resources can provide services such as frequency regulation and load balancing, contributing to the stability of the electrical grid. This study focuses on electric vehicle (EV) aggregators that participate in frequency-supporting ancillary service markets in the Nordic region [1]. However, the stochastic nature of the charging behavior and availability of electric vehicles creates significant challenges in ensuring reliable market participation, which requires strategies to manage uncertainty in the flexible supply of the aggregator.
To enable the participation of flexible yet stochastic resources in ancillary service markets, the Danish transmission system operator (TSO), Energinet, has introduced grid codes requiring a minimum 90% reliability for the full availability of ancillary service bids [2]. By explicitly allowing up to 10% failure, meaning partial or complete unavailability of the bid, this requirement encourages participation from resources with uncertain consumption or production baselines. As a result, it increases the supply of ancillary services, potentially reduces service prices, and reduces the TSO’s procurement costs. However, accommodating uncertainty in service availability during activation events may compromise grid security.
This paper examines how aggregators can leverage this requirement. The analysis of its implications for grid security and the additional service volumes the TSO may need to procure to maintain a desired confidence level is left for future work. To meet the 90% reliability requirement, aggregators must apply optimization techniques that incorporate probabilistic constraints to manage uncertainty while ensuring compliance with the TSO regulations. Chance-constrained optimization naturally lends itself to this task.
We focus on the bidding strategy of an EV aggregator that uses the 90% reliability requirement to place hourly reserve capacity bids in the Nordic ancillary service markets. Without loss of generality, we consider the market for frequency containment reserve for disturbances (FCR-D). Bidding decisions are made on the day-ahead stage, where future EV consumption remains uncertain. The key challenge is modeling these uncertainties while ensuring compliance with the TSO requirement. Furthermore, we explore how the aggregators strategy would change if the TSO adjusts the reliability threshold.
As our main methodological contribution, we develop a chance-constrained optimization model using concepts from extreme value theory and propose an analytical reformulation. This contrasts with the widely used sample-based approach, which requires an increasing number of samples to capture the underlying uncertainty.
I-B Literature review
Energinet has taken a leading role among European TSOs by implementing grid codes that define the requirements for stochastic, flexible resources providing ancillary services. Research on bidding strategies for aggregators of flexible resources, constrained by these grid codes, has explored various techniques to manage the uncertainties involved in service delivery. These techniques include robust optimization [3], distributionally robust optimization [4], and optimization methods incorporating both chance constraints and conditional value-at-risk constraints [5], often reformulated using historical samples.
A common approach to solving optimization problems with uncertainty is scenario-based methods, where uncertainties are represented by a finite set of scenarios sampled from the underlying probability distribution or empirical data [6]. While straightforward, these methods often require a large number of scenarios to accurately capture the underlying uncertainty, leading to increased computational costs [7, 8]. An alternative approach involves approximation techniques or analytical reformulations for specific types of probability distributions. These methods convert optimization problems with probabilistic constraints into analytically tractable forms, enabling more efficient solution methods [9, 10, 11, 12, 13].
A relatively underexplored but promising area in optimization under uncertainty is the analysis of the “tail” of the data, which focuses on the extreme events where failures are more likely to occur. While traditional approaches often use tail behavior or extreme values for post hoc risk assessments, there is growing interest in embedding this analysis directly into the optimization process, particularly in energy systems and resource aggregation. This gap presents an opportunity to integrate extreme value theory into chance-constrained optimization models. By doing so, decision-making under extreme uncertainties can be enhanced without the need for a large number of samples, as required by conventional scenario-based approaches. In this direction, [14] combines large deviation theory with convex analysis and bilevel optimization to derive tractable formulations that can be solved using standard solvers, assuming Gaussian mixture distributions.
@@ I-B Literature review
To the best of our knowledge, this paper is the first to propose a framework that integrates ideas from extreme value theory into chance-constrained optimization for optimizing the bidding strategies of aggregators of stochastic flexible resources in ancillary service markets, while also incorporating and analyzing the impact of real-world market regulations.
I-C Contributions
This study develops a chance-constrained optimization model for an EV aggregator in Nordic ancillary service markets, accounting for the stochastic consumption baseline of EV fleets. By applying concepts from extreme value theory to analyze the tail behavior of EV charging patterns, the model focuses on rare, high-impact events that could lead to constraint violations and, consequently, unsuccessful flexibility service delivery. This approach improves risk characterization and ensures the model effectively handles worst-case scenarios. An analytical solution is provided that is less computationally demanding compared to commonly used sample-based methods. Additionally, the study explores how varying failure thresholds set by Energinet impact the ancillary service market participation of the EV aggregator, offering practical insights for both aggregators and the TSO on risk assessment and supply of ancillary services.
I-D Paper organization
@@
The remainder of this paper is organized as follows: Section II provides the necessary preliminaries for understanding the subsequent sections. Section III presents the formulation of the chance-constrained optimization problem. Section IV outlines the methodology for the analytical reformulation of the problem, specified in Section V. Section VI discusses a common sample-based method for comparison purposes. Numerical results are presented in Section VII. Finally, Section VIII concludes the paper. For readers interested in the case study to which the methods are applied, further details are provided in Appendix A.
II Preliminaries
@@
II-A FCR-D market
Frequency containment reserve (FCR) is a category of ancillary services that helps balance short-term deviations in power and maintain the frequency stability of the electrical grid. In the Nordic grid, known for its low inertia, relatively small system size, and high share of stochastic renewable energy, FCR is divided into multiple services. These include FCR for normal operation (FCR-N), which is activated when the frequency deviates between 49.9 Hz and 50.1 Hz, and FCR for disturbances (FCR-D), which is activated when frequency deviations exceed the normal operational range. Specifically, FCR-D up services are triggered when the frequency drops below 49.9 Hz, while FCR-D down services are activated when the frequency rises above 50.1 Hz. To stabilize the grid, FCR-N and FCR-D service providers adjust their power production or consumption levels in response to frequency fluctuations. FCR-N is a symmetrical product, with equal bids made in both directions, whereas FCR-D is asymmetrical and split into two separate products: upward and downward [2]. FCR-N and FCR-D are procured in two distinct markets, each tailored to their specific operational roles. Without loss of generality, this paper focuses on the FCR-D market, as asymmetrical bidding for upward and downward services is well suited for EV aggregators and their battery state-of-charge constraints.
II-B The P90 requirement
To maintain grid stability, the provision of ancillary services must be highly reliable. However, imposing a 100% reliability requirement for ancillary service bids would limit the participation of many non-conventional technologies, such as wind power or aggregated flexible loads like EVs. To facilitate the participation of stochastic flexible resources in the FCR markets, including FCR-D, which is the focus of this paper, Energinet has introduced the P90 requirement, as outlined in [2] and [5].
The P90 requirement ensures that flexibility providers, such as EV aggregators, can participate in the FCR markets, provided that their reserved capacity bid is fully realized at least 90% of the time. To be eligible to bid into the markets, an aggregator must complete Energinet’s prequalification process [2] and conduct a prognosis analysis of its ability to meet future demand, which is verified by Energinet. Once qualified by the TSO, a flexibility provider can participate in the relevant ancillary service markets. From that point, Energinet continuously performs ex-post checks to verify whether a reserved capacity bid by the participant is truly available, regardless of whether the reserve is activated. This serves as an additional security measure for supply. For example, an upwards reserve bid of 10 kW is considered unavailable if the actual consumption realization of an EV aggregator is only 8 kW. Any partial or full unavailability of a bid counts as a reserve shortfall. Energinet conducts these checks at appropriate time intervals, such as every quarter. If a provider violates the P90 requirement, it will be excluded from further market participation and must re-apply for qualification.
This requirement can be expressed mathematically using chance constraints of the form $\mathbb{P}(\cdot)\geq 1-\varepsilon$, where $\varepsilon=0.1$ represents the maximum 10% allowance for failure in successfully delivering the service.
II-C The LER requirement
In addition to the P90 requirement, Energinet has introduced an additional requirement for resources with limited energy reservoirs (LER), such as batteries and EV aggregators. The primary objective of the LER requirement is to ensure that resources with limited energy capacity can maintain energy availability despite their constraints. This includes adhering to energy management protocols and ensuring the ability to sustain continuous activation for a specified minimum duration when needed. The definition of this requirement is outlined in [2] and [5].
The specifics of the LER requirement differ between the FCR-N and FCR-D services. This paper focuses on FCR-D services, where the requirement states that an LER unit bidding in one direction of the FCR-D market (either up or down) must also ensure that at least 20% of the reserve bid is available in the opposite direction. For EV aggregators, this LER requirement applies only in the downwards direction, since activation in this direction (i.e., increasing power consumption) is constrained by the battery energy storage capacity of EVs. In contrast, activation in the upwards direction (i.e., reducing power consumption) is not constrained by battery capacity and is treated similarly to other non-LER flexible loads. Additionally, the requirement mandates that the service must be available for at least 20 continuous minutes once activated.
@@ II-C The LER requirement
where $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$ are random variables representing the available upwards, downwards, and energy flexibility, respectively. Downwards flexibility, $R^{\downarrow}$, refers to the potential for increasing consumption of the grid-connected EV fleet, while upwards flexibility, $R^{\uparrow}$, refers to the potential for reducing consumption. In addition, energy flexibility, $R^{\rm{E_{20}}}$, is determined based on the current state of charge of the entire EV fleet and the maximum power that can be applied to the fleet over the next 20 minutes without exceeding battery capacity.
The realizations of these random variables are denoted as $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$. A detailed explanation of how these realizations are defined can be found in Appendix A. For a quicker overview, Fig. 1 provides graphical representations of how the flexibilities are estimated. Specifically, the realizations $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$ are defined as the extreme values of flexibility within the hour, which in this study corresponds to the minimum available flexibility during the hour. This approach differs from [5], where similar sets are defined using minute-level realizations. By considering the minimum value over the entire hour, the computational burden is reduced, as the number of data points decreases by a factor of 60.
II-D Case study
Our case study utilizes real-world data comprising charging profiles from 1400 residential EV charging boxes in Denmark. The measurements were recorded from March 24, 2022, to March 21, 2023, with an average time step of 2.84 minutes. To achieve a uniform resolution, the data has been interpolated to a one-minute granularity. It is assumed that each charging box is associated with a single EV. The historical EV consumption levels serve as the baseline for estimating each flexibility. Further details are provided in Appendix A.
III Chance-constrained bidding model
Figure 2: Scatter plots illustrating the estimated upwards (left), downwards (middle), and energy (right) flexibilities for an arbitrary hour across each day of the year. The blue dots represent the in-sample data used for optimization in a given run, while the pink dots are used for the ex-post out-of-sample validation in the same run. The dashed grey line indicates the 10th percentile of the blue dots. In line with extreme value theory, only the samples below this percentile are used for estimation in the analytical method.
@@ III Chance-constrained bidding model
In line with the P90 and LER requirements, we adopt a joint chance-constrained optimization model similar to the one proposed in [5]:
@@
where the objective function (2a) maximizes the combined reserve capacity bids into the FCR-D up and down markets. Additionally, (2b) defines the joint chance constraint that incorporates both the LER and P90 requirements. Specifically, the chance constraint $\mathbb{P}(\cdot)\geq 1-\varepsilon$ ensures that the probability of jointly satisfying the constraints within the corresponding hour is at least $1-\varepsilon$. In accordance with the P90 requirement, $\varepsilon=0.1$. Note that the bids are submitted on an hourly basis, and the proposed model accounts only for the reservation of flexibility capacity. As a result, the rebound effect is not considered, meaning there is no inter-hour dependency in the bidding scheme. Future work could address this by incorporating activation data.
@@
Chance-constrained programs are generally computationally intractable. However, if the underlying probability distribution satisfies certain properties, analytical reformulations of the chance constraints can be derived. For example, analytical methods like the Bonferroni approximation [15] transform probabilistic constraints into a series of more tractable constraints, simplifying the problem at the potential cost of precision. Alternatively, sample-based methods represent uncertainty through a set of likely scenarios, allowing for optimization that accounts for a range of possible outcomes. These approaches enable EV aggregators to optimize their bidding strategies while managing the stochastic nature of their resources within the operational constraints of ancillary service markets.
@@
In Fig. 2, scatter plots of the in-sample and out-of-sample data used in our case study are shown. In-sample data refers to the set used in the optimization problem to determine optimal bids, while out-of-sample data is used ex-post for bid evaluation. Extreme value theory [16] focuses on the behavior of extreme values (maxima or minima) within a dataset, particularly as the sample size increases. We are particularly interested in the extreme values, which in our case correspond to those below the dotted line in Fig. 2, representing the empirical 10th percentile of the in-sample data. To avoid bias, we solve the model in multiple runs, each with different sets of in-sample and out-of-sample data, while keeping the number of samples in each set unchanged.
@@
To derive an analytical solution to (2), we aim to fit a distribution to the tail data (samples below the empirical 10th percentile) that facilitates an analytical reformulation of the chance constraint. Among the various options for modeling tail data, we tested both the Weibull and Generalized Pareto distributions. We observed better out-of-sample performance with the former, and thus continue our analysis using the Weibull distribution.
@@
It is worth noting that in the literature on extreme value theory, the generalized extreme value distribution is more commonly used, as it emerges as the limiting distribution of the maximum of $n$ independent and identically distributed random variables when $n$ grows large. However, the convergence to this limiting distribution may be poor. Additionally, the generalized extreme value distribution introduces extra parameters that require tuning, which may also demand more samples.
IV Distribution fitting and validation

### v4

@@ Abstract
To enable the participation of stochastic distributed energy resources in ancillary service markets, the Danish transmission system operator, Energinet, mandates that flexibility providers satisfy a minimum 90% reliability requirement for reserve bids. This paper examines the bidding strategy of an electric vehicle aggregator under this regulation and develops a chance-constrained optimization model. In contrast to conventional sample-based approaches that demand large datasets to capture uncertainty, we propose an analytical reformulation that leverages extreme value theory to characterize the tail behavior of flexibility distributions. A case study with real-world charging data from 1400 residential electric vehicles in Denmark demonstrates that the analytical solution improves out-of-sample reliability, reducing bid violation rates by up to 8% relative to a sample-based benchmark. The method is also computationally more efficient, solving optimization problems up to 4.8 times faster while requiring substantially fewer samples to ensure compliance. Moreover, the proposed approach enables the construction of feasible bids with reliability levels as high as 99.95%, which would otherwise require prohibitively large scenario sets under the sample-based method. Beyond its computational and reliability advantages, the framework also provides actionable insights into how reliability thresholds influence aggregator bidding behavior and market participation. This study establishes a regulation-compliant, tractable, and risk-aware bidding methodology for stochastic flexibility aggregators, enhancing both market efficiency and power system security.
keywords:
Ancillary service markets, electric vehicle aggregation, chance-constrained optimization, extreme value theory, uncertainty modeling, reserve capacity bidding
[…]
Acknowledgment The authors would like to thank Spirii for providing the data, as well as Gustav A. Lunde and Emil V. Damm for their extensive work on data preparation, which made this case study possible. The authors would also like to thank the editor and the two anonymous reviewers for their constructive feedback, which has greatly improved the quality of this paper.
1 Introduction
1.1 Background and motivation
Flexibility aggregators pool and manage distributed energy resources such as electric vehicles (EVs), enabling them to provide ancillary services like frequency regulation and load balancing (Energinet1). By doing so, they contribute to power system stability while creating new revenue streams for consumers. However, the inherently stochastic nature of distributed resources, arising from uncertain consumption, charging, and availability patterns, makes it difficult to guarantee reliable service delivery. For system operators, ensuring that contracted reserves are available during activation events is essential to maintaining grid security.
To address this challenge, the Danish transmission system operator, Energinet, has introduced grid codes that require at least 90% reliability for ancillary service bids (Energinet). This requirement, known as the P90 requirement, is particularly relevant for EV aggregators, who must balance uncertain user behavior with regulatory compliance in order to participate in the Nordic market for ancillary services. While the requirement facilitates the market participation of stochastic flexibility providers, it also raises important methodological questions. The key challenge for aggregators is to design regulation-compliant bidding strategies that explicitly manage uncertainty without being either overly conservative or computationally intractable.
Existing optimization techniques provide partial answers but remain limited. Robust optimization guarantees feasibility under worst-case realizations, but the resulting bids are often too conservative, leading to under utilization of available flexibility. Distributionally robust optimization relaxes this conservatism by considering ambiguity sets, yet its effectiveness depends on assumptions about distributional properties that may not hold in practice. Scenario-based and sample-based approaches are widely applied due to their simplicity, but they require very large numbers of samples to approximate low-probability high-impact events, which leads to significant computational costs. More critically, these methods fail to capture extreme shortfalls that drive Energinet’s reliability checks, underscoring the need for tractable bidding models that explicit target tail risks in flexibility estimation.
1.2 Literature review
A growing body of research has examined bidding strategies for aggregators of flexible resources under uncertainty. Robust optimization approaches, such as those developed by kuhn, directly encode reliability requirements but often yield highly conservative solutions. GadeP90 propose a bidding strategy for EV aggregators using distributionally robust optimization, aiming to balance tractability and robustness, though it relies heavily on assumptions regarding the ambiguity sets. More recent studies incorporate chance constraints and conditional value-at-risk into aggregator models (GandE), yet these methods typically rely on sample approximations and therefore inherit the scalability and accuracy limitations of scenario-based approaches.
Scenario-based optimization remains the most commonly adopted method, representing uncertainty through a finite set of samples. While conceptually straightforward, it often requires hundreds or thousands of samples to capture extreme realizations, substantially increasing computational costs (BertScaling, blanchet2024efficient). Approximation techniques, such as Bonferroni-type bounds or convex reformulations (Prekopa1970, NemirovskiStewart2012), improve tractability but generally fail to capture the “tail” behavior of uncertainty distributions, where rare yet extreme shortfalls occur.
A fundamental limitation emerges across these studies: each method resolves part of the challenge, but none reconciles the trade-off between over-conservatism, restrictive assumptions, scalability, and accurately capturing tail risks. Extreme value theory offers a pathway to accurately model extreme events while preserving tractability (diebold2000pitfalls). Traditional tail and extreme value analysis focuses on post hoc risk assessment (husler2003optimization). Yet, in energy systems, where extreme deviations in supply or demand can have catastrophic impacts on system stability, there is a strong motivation to embed such analysis within the optimization process itself. This presents an opportunity to integrate concepts of extreme value theory into chance-constrained optimization, enabling decision-making under extreme uncertainties without the need for large scenario sets. In this direction, RareCCsTong combine large deviation theory with convex analysis and bilevel optimization to derive tractable formulations solvable with standard solvers, assuming Gaussian mixture models (GMMs). The inherently light-tailed structure of GMMs presents an important limitation for concrete power system applications, potentially underestimating infrequent but severe flexibility deviations. In contrast, extreme value theory is explicitly designed to model such rare events, providing a more reliable foundation for chance-constrained formulations in this context.
@@ 1.2 Literature review
To the best of our knowledge, this paper is the first to embed concepts from extreme value theory within chance-constrained optimization and demonstrate their value in practice by providing a bidding method for aggregators of stochastic flexible resources in ancillary service markets, grounded in real data to capture tail behavior, while also deriving actionable insights on the role and impact of market regulations.
1.3 Contributions
This paper addresses these gaps by developing an analytical approach to aggregator bidding that integrates extreme value theory into a chance constrained optimization framework. The main contributions are summarized as follows:
  • A chance constrained bidding model for an electric vehicle aggregator in Nordic ancillary service markets that directly incorporates Energinet’s 90% reliability requirement together with the limited energy reservoir constraint.
  • An analytical reformulation based on extreme value theory that focuses explicitly on the tail of flexibility distributions, enabling tractable solutions with far fewer samples than conventional scenario-based methods.
  • Numerical validation with real-world data from 1400 residential charging points in Denmark, showing that the extreme value theory based approach reduces out-of-sample violation rates by up to 8% and solves optimization problems up to 4.8 times faster than a sample-based benchmark.
  • A demonstration of stricter reliability compliance, where the proposed model produces feasible bids at reliability levels as high as 99.95%, which would require prohibitively large scenario sets under a sample-based approach.
  • Practical insights for aggregators and system operators, showing how changes in reliability thresholds influence bidding strategies, revenue opportunities, and market participation.
Together, these contributions establish a novel, regulation-compliant, and computationally efficient framework for stochastic flexibility bidding, enhancing both market efficiency and power system security.
1.4 Paper organization
@@
The remainder of this paper is structured as follows. Section 2 introduces the necessary preliminaries for understanding the subsequent sections. Section 3 presents the formulation of the chance-constrained optimization problem, while Section 4 outlines the methodology for its analytical reformulation, detailed further in Section 5. Section 6 describes a commonly used sample-based approach for comparison. Numerical results are reported in Section 7, and Section 8 concludes the paper. Additional details on the case study and the distribution fitting procedure are provided in Appendices A and B.
2 Preliminaries
2.1 FCR-D market
Frequency containment reserve (FCR) is a category of ancillary services that helps balance short-term deviations in power and maintain the frequency stability of the electrical grid. In the Nordic grid, known for its low inertia, relatively small system size, and high share of stochastic renewable energy, FCR is divided into multiple services. These include FCR for normal operation (FCR-N), which is activated when the frequency deviates between 49.9 Hz and 50.1 Hz, and FCR for disturbances (FCR-D), which is activated when frequency deviations exceed the normal operational range. Specifically, FCR-D up service is triggered when the frequency drops below 49.9 Hz, while FCR-D down service is activated when the frequency rises above 50.1 Hz. To stabilize the grid, FCR-N and FCR-D service providers adjust their power production or consumption levels in response to frequency fluctuations. FCR-N is a symmetrical product, with equal bids made in both directions, whereas FCR-D is asymmetrical and split into two separate products: upward and downward (Energinet). FCR-N and FCR-D are procured in separate day-ahead markets, each tailored to their specific operational roles. Without loss of generality, this paper focuses on the FCR-D market, since its asymmetric bidding structure for upward and downward services aligns well with the characteristics of EV aggregators and their battery state-of-charge constraints.
2.2 The P90 requirement
To maintain grid stability, the provision of ancillary services must be highly reliable. However, imposing a 100% reliability requirement for ancillary service bids would limit the participation of many non-conventional technologies, such as wind power or aggregated flexible loads like EVs. To facilitate the participation of stochastic flexible resources in ancillary service markets, including FCR-D, which is the focus of this paper, Energinet has introduced the P90 requirement, as outlined in Energinet (Energinet) and GandE (GandE).
The P90 requirement ensures that flexibility providers, such as EV aggregators, can participate in ancillary service markets, provided that their reserved capacity bid is fully realized at least 90% of the time. To be eligible to bid into the markets, an aggregator must complete Energinet’s prequalification process (Energinet) and conduct a prognosis analysis of its ability to meet future demand, which is verified by Energinet. Once qualified by Energinet, a flexibility provider can participate in the relevant ancillary service markets. From that point, Energinet continuously performs ex-post checks to verify whether a reserved capacity bid by the participant is truly available, regardless of whether the reserve is activated. This serves as an additional security measure for supply. For example, an FCR-D up-reserve bid of 10 kW becomes unavailable if the actual consumption of an EV aggregator is only 8 kW, since the aggregator cannot reduce its consumption by the full 10 kW. Any partial or full unavailability of a bid counts as a reserve shortfall. Energinet conducts these checks at appropriate time intervals, such as every yearly quarter. If a provider violates the P90 requirement, meaning that the frequency of non-zero shortfalls exceeds 10% of the time, it will be excluded from further market participation and must re-apply for qualification.
In the context of modelling such requirements mathematically, this requirement can be expressed using chance constraints of the form $\mathbb{P}(\cdot)\geq 1-\varepsilon$, where $\varepsilon=0.1$ represents the maximum 10% allowance for failure in successfully delivering the service.
2.3 The LER requirement
In addition to the P90 requirement, Energinet has introduced an additional requirement for resources with limited energy reservoirs (LER), such as batteries and EV aggregators. The primary objective of the LER requirement is to ensure that resources with limited energy capacity can maintain energy availability despite their constraints. This includes adhering to energy management protocols and ensuring the ability to sustain continuous activation for a specified minimum duration when needed. The definition of this requirement is outlined in Energinet (Energinet) and GandE (GandE).
The specifics of the LER requirement differ between FCR-N and FCR-D services. This paper focuses on FCR-D, where an LER unit bidding in one direction of the market (either up or down) must also ensure that at least 20% of its capacity is available in the opposite direction. Mathematically, this is captured in (1a), which requires reserving 20% of the available upward flexibility for downward activation. For EV aggregators, this LER requirement applies only in the upward direction, because downward activation (i.e., increasing power consumption) is limited by the battery storage capacity of the EV. In contrast, upward activation (i.e., reducing power consumption) is not constrained by battery capacity since the power supply to a charging EV can always be turned off and is treated like other non-LER flexible loads. The requirement further specifies that the service must remain available for at least 20 continuous minutes once activated.
(a) Downwards and upwards flexibility for an aggregated EV fleet.
(b) Energy flexibility for a single EV and for an aggregated fleet.
@@ 2.3 The LER requirement
where $R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}$} are random variables representing the available upward, downward, and energy flexibility, respectively. Specifically, upward flexibility $R^{\uparrow}$ in (1a) refers to the potential for reducing consumption of the grid-connected EV fleet and is subject to the LER requirement described above. Downward flexibility $R^{\downarrow}$ in (1b) refers to the potential for increasing consumption. Energy flexibility $R^{\rm{E_{20}}}$ in (1c) is determined by the current state of charge of the entire EV fleet and the maximum power that can be applied over the next 20 minutes without exceeding battery capacity.
The realizations of these random variables are denoted as $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$. A detailed explanation of how these realizations are defined can be found in Appendix A. For a quicker overview, Figure 1 provides graphical representations of how the flexibilities are estimated. Specifically, the realizations $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$ are defined as the extreme values of flexibility within the hour, which in this study corresponds to the minimum available flexibility during the hour. In contrast to GandE, who define similar sets using minute-level realizations, our hourly approach reduces computational burden by decreasing the number of data points by a factor of 60 while also mitigating temporal dependencies in the time series.
3 Chance-constrained bidding model
@@
In line with the P90 and LER requirements, we adopt a joint chance-constrained optimization model similar to the one proposed in GandE (GandE):
@@
The objective function in (2a) maximizes the combined reserve capacity bids in the FCR-D up and down markets. Constraint (2b) defines the joint chance constraint, which incorporates both the LER and P90 requirements. Specifically, the chance constraint $\mathbb{P}(\cdot)\geq 1-\varepsilon$ ensures that the probability of jointly satisfying these constraints within a given hour is at least $1-\varepsilon$. In accordance with the P90 requirement, $\varepsilon=0.1$. The bids are submitted on an hourly basis (though in Denmark this will be changed to 15-minute intervals), and the proposed model accounts only for the reservation of flexibility capacity. Consequently, the rebound effect is not considered, implying no inter-temporal dependency within the hour in the bidding scheme. Future work could address this limitation by incorporating activation data.
@@
Chance-constrained programs are generally computationally intractable. However, if the underlying probability distribution satisfies certain properties, the chance constraints can be reformulated analytically. For instance, methods such as the Bonferroni approximation (BonferroniCorr) replace probabilistic constraints with a series of more tractable ones, simplifying the problem by evaluating only the marginal distributions of the random variables, albeit at the cost of precision. This means that any correlation structure among the random variables $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$ in the joint chance constraint is disregarded. Alternatively, sample-based methods capture uncertainty through a set of representative scenarios, enabling optimization that explicitly accounts for a range of possible outcomes. Both approaches allow EV aggregators to design bidding strategies that manage the stochastic nature of their resources while respecting the operational requirements of ancillary service markets.
Figure 2: Scatter plots illustrating the estimated upward (top), downward (middle), and energy (bottom) flexibilities for a representative hour across all days of the year. Blue samples represent the in-sample data used for optimization in a given run, while pink samples correspond to the ex-post out-of-sample validation in the same run. The dashed line marks the empirical 10th percentile of the blue dots. Consistent with extreme value theory, only the blue samples below this percentile are used for estimation in the analytical method.
@@
Figure 2 presents scatter plots of the in-sample and out-of-sample data used in our case study. The in-sample data (blue samples) are used in the optimization problem to determine the optimal bids, while the out-of-sample data (pink samples) are used ex-post to evaluate these bids. Extreme value theory (EVTCastillo) studies the behavior of extreme values (maxima or minima) in a dataset, particularly as the sample size increases. In this study, the extreme values of interest correspond to the blue samples below the dashed line, which represents the empirical 10th percentile of the in-sample data. To reduce bias, the model is solved 10 times, each with a different set of in-sample and out-of-sample data, while keeping the number of samples in each set constant.
@@
To derive an analytical reformulation of the optimization problem (2), we fit a distribution to the tail data (blue samples below the empirical 10th percentile) that enables an analytical representation of the chance constraint. Among the options considered for modeling the tail, we tested both the Weibull and Generalized Pareto distributions. The Weibull distribution was preferred because it is comparatively more straightforward to fit and tune, yields a lower negative log-likelihood when fitted via maximum likelihood, and demonstrates better out-of-sample performance. Consequently, our analysis is based on the Weibull distribution.
@@
It is worth noting that in the extreme value theory literature, the generalized extreme value distribution is more commonly used, as it arises as the limiting distribution of the maximum of $n$ independent and identically distributed random variables as $n$ becomes large. However, convergence to this limiting distribution can be slow. Moreover, the generalized extreme value distribution introduces additional parameters that must be estimated, often requiring a larger number of samples.
4 Distribution fitting and validation